# An Efficient Sparse Kernel Generator for O(3)-Equivariant Deep Networks

### v1

@@ 4.2 Roofline Analysis
B 12.3 - 11.35
Table 2: Arithmetic intensity (AI) and throughput for our kernels vs. cuEquivariance (cuE) for varied problem configurations, FP32 forward and backward passes. cuE does not run the backward pass on these configurations due to a documented issue.
@@
Figure 7: Roofline analysis of our kernels on the input configurations from Table 2. Our kernel performance closely tracks the slope of the global memory roofline, whereas cuE plateaus early in the memory-bandwidth limited region of the plot.
@@ 4.3 Kernel Fusion Benchmarks
The“fast-fused-det” and “fast-fused-atomic” implementations use the memory-efficient formulations from Section 3.5, offering dramatically faster performance as a result. Our deterministic fused algorithm offers up to 4.5x speedup over cuEquivariance for the forward pass and 3.5x on the backward pass. Our poorest performance occurs with the atomic algorithm in FP64 precision for the backward pass, where we still see 1.3x-1.4x speedup over cuEquivariance.
@@ 4.4 Acceleration of MACE
Figure 9 (left) compares the rate of molecular dynamics simulation among the different kernel providers. We benchmarked cuE with the optimal data layout for its irreps and included optimizations for symmetric tensor contraction, linear combination layers, and graph convolution. In FP32 precision, we provide 5.3x speeedup over e3nn and 1.5x over cuEquivariance. In FP64 precision, we offer 6.6x speedup over e3nn and 1.43x over cuEquivariance. Because we do not materialize the intermediate tensor product outputs for each edge, our fused kernels reduce the model memory footprint by roughly 158x. Figure 9 (right) breaks down the device runtime spent in various kernels traced by the Pytorch profiler. The sum of kernel runtimes in the right figure differs from the times measured in the left figure due to device idle time. Notice that our speedup comes entirely from accelerating the equivariant graph convolution; the CTP kernel and scatter-sum runtime is 4x faster in our implementation compared to cuE in FP32 precision. By contrast, cuEquivariance also accelerates a structured linear layer and a “symmetric contraction” kernel specific to MACE (reducing the runtime counted in “Other” category), and we leave these kernels to future work. Our runtime in the “Other” category can be reduced drastically through torch.compile(), but we did not include this optimization in fairness to cuE (lack of support) and e3nn (memory constraints).
@@ 5 Conclusions and Further Work
  • Integration into new models: Our software remains accessible to newcomers while delivering the high performance required for massive inference workloads. In conjuction with domain experts, we hope to apply our library to train larger, more expressive equivariant deep neural networks that deliver performance at scale.

### v3

@@ 4.2 Roofline Analysis
B 12.3 1.56 11.30
Table 2: Arithmetic intensity (AI) and throughput for our kernels vs. cuEquivariance (cuE) for varied problem configurations, FP32 forward and backward passes.
@@
Figure 7: Roofline analysis of our kernels on the input configurations from Table 2, batch size 200k. Our kernel performance closely tracks the slope of the global memory roofline, whereas cuE plateaus early in the memory-bandwidth limited region of the plot.
@@ 4.3 Kernel Fusion Benchmarks
The“fast-fused-det” and “fast-fused-atomic” implementations use the memory-efficient formulations from Section 3.5, offering dramatically faster performance as a result. Our deterministic fused algorithm offers up to 4.8x speedup over cuEquivariance for the forward pass and 3.5x on the backward pass. Our poorest performance occurs with the atomic algorithm in FP64 precision for the backward pass, where we still see 1.3x-1.4x speedup over cuEquivariance.
@@ 4.4 Acceleration of MACE
Figure 9 (left) compares the rate of molecular dynamics simulation among the different kernel providers. We benchmarked cuE with the optimal data layout for its irreps and included optimizations for symmetric tensor contraction, linear combination layers, and graph convolution. In FP32 precision, we provide 5.3x speeedup over e3nn and 1.5x over cuEquivariance. In FP64 precision, we offer 6.6x speedup over e3nn and 1.43x over cuEquivariance. Because we do not materialize the intermediate tensor product outputs for each edge, our fused kernels reduce the model memory footprint as well.
Figure 9 (right) breaks down the device runtime spent in various kernels traced by the PyTorch profiler. The sum of kernel runtimes in the right figure differs from the times measured in the left figure due to device idle time. Notice that our speedup comes entirely from accelerating the equivariant graph convolution; the CTP kernel and scatter-sum runtime is 4x faster in our implementation compared to cuE in FP32 precision.
By contrast, cuEquivariance also accelerates a symmetric contraction kernel specific to MACE (reducing the runtime counted in “Other” category), and we leave this kernel to future work. Our runtime in the “Other” category can be reduced through torch.compile(), but we did not include this optimization in fairness to cuE (compilation erros in MACE) and e3nn (memory constraints).
@@ 5 Conclusions and Further Work
  • Integration into new models: Our software remains accessible to newcomers while delivering the high performance required for massive inference workloads. In conjunction with domain experts, we hope to apply our library to train larger, more expressive equivariant deep neural networks.